OpenStudy (anonymous):

SOMEONE PLEASE HELP ASAP. Find the equation of a circle with a center at (7, 2) and a point on the circle at (-2, -5).

OpenStudy (anonymous):

so the formula you want to use is (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

what do I plug in though?

OpenStudy (anonymous):

Plug in your point for x and y

OpenStudy (anonymous):

what is h and k?

OpenStudy (anonymous):

wait, before you find the equation... we need to find the radius

OpenStudy (anonymous):

since you are given two points

OpenStudy (anonymous):

d = sqrt(x^2-x^1)^2 + (y^2-y^1)^2

OpenStudy (anonymous):

tell me what you get for the radius

OpenStudy (anonymous):

once you found the radius, you can plug in the radius into the equation for the circle

OpenStudy (anonymous):

could you possibly plug in the numbers cause im confused

OpenStudy (anonymous):

idk how to get the radius though

OpenStudy (anonymous):

d = sqrt(-2-7)^2 + (-5-2)^2

OpenStudy (anonymous):

130 right?

OpenStudy (anonymous):

yes, now plug in 130 into the other formula